# A Comprehensive Survey on Instruction Following

### v7

@@ Abstract
Task semantics can be expressed by a set of input-output examples or a piece of textual instruction. Conventional machine learning approaches for natural language processing (NLP) mainly rely on the availability of large-scale sets of task-specific examples. Two issues arise: first, collecting task-specific labeled examples does not apply to scenarios where tasks may be too complicated or costly to annotate, or the system is required to handle a new task immediately; second, this is not user-friendly since end-users are probably more willing to provide task description rather than a set of examples before using the system. Therefore, the community is paying increasing interest in a new supervision-seeking paradigm for NLP: learning to follow task instructions, i.e., instruction following. Despite its impressive progress, there are some common issues that the community struggles with. This survey paper tries to summarize and provide insights to the current research on instruction following, particularly, by answering the following questions: (i) What is task instruction, and what instruction types exist? (ii) How to model instructions? (iii) What are popular instruction following datasets and evaluation metrics? (iv) What factors influence and explain the instructions’ performance? (v) What challenges remain in instruction following? To our knowledge, this is the first comprehensive survey about instruction following. [^1] The curated paper list can be found at:  https://github.com/RenzeLou/awesome-instruction-learning
@@ 1 Introduction
One goal of AI is to build a system that can universally understand and solve new tasks. Labeled examples (Figure 1 (a)), as the mainstream task representation, are costly to obtain at scale or even do not exist in some cases. Then, is there any other task representation that can contribute to task comprehension? Textual instructions provide another dimension of supervision for expressing the task semantics, which often contains more abstract and comprehensive knowledge of the target task than individual labeled examples. As shown in Figure 1 (b), with the availability of task instructions, systems can be quickly built to handle new tasks. Such efficiency is highly desirable in real-world applications, especially when task-specific annotations are scarce. More importantly, instruction following leans toward human intelligence in terms of learning new tasks—a little child can well solve a new mathematical task by learning from its instruction and a few examples Fennema et al. (1996); Carpenter et al. (1996). As a result, this new learning paradigm has recently attracted the main attention of the machine learning and NLP communities Wang et al. (2022d); Longpre et al. (2023).
@@ 2.1 Instruction Following
As illustrated in Figure 1, unlike traditional example-driven supervised learning, the essence of instruction following is to train the LLMs to understand various instructions and produce the corresponding responses. Since this capacity can be extended to any unseen downstream tasks, instruction following has become an efficient learning paradigm for solving few/zero-shot tasks (Radford et al., 2019; Schick and Schütze, 2021c; Yin et al., 2022; Li et al., 2023a; Gupta et al., 2023, inter alia). However, the performance of instruction following highly relies on both model and task scale: a larger LLM (or pretraining with more tokens) tuned on more diverse tasks can achieve significantly better few/zero-shot performances on the downstream tasks (Chung et al., 2022; Iyer et al., 2022; Wang et al., 2023b, inter alia). As scaling model size is unrealistic for most of us, numerous recent studies worked on collecting high-quality instruction-tuning datasets, either employing human labors Khashabi et al. (2020); Ye et al. (2021); Sanh et al. (2022); Wang et al. (2022d); Longpre et al. (2023); Conover et al. (2023); Köpf et al. (2023) or distilling supervision from the powerful LLMs Wang et al. (2022c); Honovich et al. (2022a); Taori et al. (2023); Peng et al. (2023); Xu et al. (2023b); Geng et al. (2023); Chiang et al. (2023); Xu et al. (2023a); Köksal et al. (2023); Kim et al. (2023); Ding et al. (2023); Yin et al. (2023), e.g., utilizing ChatGPT or GPT-4 to develop creative task instructions OpenAI (2022, 2023).
@@ 2.2 Surveys on In-context Instructions
Several existing surveys Dong et al. (2023); Huang and Chang (2022); Qiao et al. (2022); Yu et al. (2023a) share similar motivations with us while focusing on merely some sub-area of instruction following, such as prompt, few-shot demonstrations, chain-of-thoughts reasoning, etc. For example, Liu et al. (2023a) provided a comprehensive overview of prompt learning and LLMs, where the prompt can be regarded as one specific type of textual instruction (as categorized in § 4). Some other studies surveying “soft instruction”, namely parameter-efficient fine-tuning methods Lialin et al. (2023), also differ from our scope of “textual instruction”. To the best of our knowledge, this is the first work that provides a comprehensive and high-level story of instruction following.
@@ 3 Preliminary
• Template (T): a textual template that either tries to express task intent or is used for bridging X and Y. [^2] A plain template connecting X and Y, e.g., “The input is [normal-$\ldots$] The output is [normal-$\ldots$]”, no task-specific semantics. T may not be an instruction yet.
@@ 4 What is Task Instruction?—A Unified Perspective from Indirect Supervision
Task
TE premise (i.e., input text) TE hypothesis (i.e., instructions Y)
@@
[Donald Trump]${}_{ent}$ served as the 45th president of the United States from 2017 to 2021. ($\checkmark$) Donald Trump is a politician
@@
[Donald Trump]${}_{ent1}$ served as the 45th president of the [United States]${}_{ent2}$ from 2017 to 2021. ($\checkmark$) Donald Trump is citizen of United States
@@
In [1997]${}_{time}$, the [company]${}_{sub}$ [hired]${}_{trigger}$ [John D. Idol]${}_{obj}$ to take over Bill Thang as the new chief executive. ($\checkmark$) John D. Idol was hired.
@@
Salesforce and Slack Technologies have [entered]${}_{event1}$ into a definitive agreement] under which Salesforce will [acquire]${}_{event2}$ Slack. ($\checkmark$) Salesforce acquires Slack after it enters into the agreement with Slack Tech.
@@
Task
Input X Template T (cloze question)
Answer
Output Y
Sentiment Analysis I would like to buy it again. [X] The product is    .
Great
@@
$\ldots$
Positive
Entity Tagging [Donald Trump]${}_{ent}$ served as the 45th president of the United States from 2017 to 2021. The entity in [X] is a    class?
Politician
@@
$\ldots$
People
[…]
Executive
@@
$\ldots$
President
[…]
Indeed
@@
$\ldots$
Yes
Translation Donald Trump served as the 45th president of the United States from 2017 to 2021. Translate [X] to French:
été président …
@@ 4.1 Three Types of Instructions
The advantages of NLI-oriented Instructions learning are four-fold: (i) it keeps the label semantics and makes it possible to encode the input-output relations; (ii) it unifies various classification problems into an NLI task; (iii) by making use of the indirect supervision from existing NLI datasets, a model trained on NLI tasks is expected to work on other tasks in a zero-shot manner; (iv) it extends the original close-set indices classification problem into an open-domain label recognition paradigm. Therefore, it has been widely used in a variety of few/zero-shot classification tasks Xu et al. (2022a), such as classifying topics Yin et al. (2019), sentiments Zhong et al. (2021), stances Xu et al. (2022c), entity types Li et al. (2022a), entity relations Murty et al. (2020); Xia et al. (2021); Sainz et al. (2021, 2022), etc.
Task
Input X Template T + Few-shot Demonstrations
Output Y
@@
 Input: [X]  Output:
Positive
@@
 Input: [X]  Output:
Donald Trump
@@
As shown in Figure 2 (b) and Table 2, the prompt is a representative of the LLM-oriented Instructions, which is usually a brief utterance prepended with the task input (prefix prompt), or a cloze-question template (cloze prompt). It is basically designed for querying the intermedia responses (that can be further converted into the final outputs) from the LLM. Since the prompted input conforms to the pre-training objectives of LLM (e.g., the cloze-style input satisfies the masked language modeling objective Kenton and Toutanova (2019)), it helps get rid of the reliance on the traditional supervised fine-tuning and greatly alleviates the cost of human annotations. Thus, prompt learning achieved impressive results on a multitude of previous few/zero-shot NLP tasks, like question answering Radford et al. (2019); Lin et al. (2021), machine translation Li et al. (2022c), sentiment analysis Wu and Shi (2022), textual entailment Schick and Schütze (2021a, b), entity recognition Cui et al. (2021); Wang et al. (2022a), etc.
Despite the excellent performance of prompt techniques, there are still two obvious shortcomings with LLM-oriented Instructions in real-world applications. (i) Not User-Friendly. As the prompt is crafted for serving LLMs, it is encouraged to design the prompt in a “model’s language” (e.g., model-preferred incoherent words or internal embedding). However, this LLM-oriented style is hard to be understood by users and often violates human intuitions Gao et al. (2021); Li and Liang (2021); Qin and Eisner (2021); Khashabi et al. (2022). Meanwhile, the performance of prompts highly depends on the laborious prompt engineering Bach et al. (2022), but most end-users are not LLM experts and usually lack sufficient knowledge to tune an effective prompt. (ii) Applications Constraints. The prompt is usually short and simplistic, whereas many tasks cannot be effectively formulated with solely a brief prompt, making prompt hard to deal with the diverse formats of real-world NLP tasks Chen et al. (2022b); Zhang et al. (2023).
@@
Human-oriented Instructions essentially denotes the instructions used for crowd-sourcing on the human-annotation platforms (e.g., Amazon MTurk). Unlike LLM-oriented Instructions, Human-oriented Instructions  (Figure 2 (c)) are usually some human-readable, descriptive, and paragraph-style information consisting of various components, such as “task title”, “category”, “definition”, and “things to avoid”, etc. (cf. Mishra et al., 2022b) Thus, Human-oriented Instructions are more user-friendly and can be ideally applied to almost any complex NLP task. Table 3 further shows some representative task examples.
Accordingly, Human-oriented Instructions have attracted much more attention in recent years (Hu et al., 2022; Gupta et al., 2022; Yin et al., 2022, inter alia). However, due to the complex nature, Human-oriented Instructions are more challenging to encode by vanilla LLMs. For example, off-the-shelf GPT-2 was found to work poorly on following MTurk instructions Wolf et al. (2019); Efrat and Levy (2020). To tame the LLMs better understand the Human-oriented Instructions, follow-up works began to collect large-scale instruction datasets Mishra et al. (2022b); Wang et al. (2022d). All the previous results showed that, after fine-tuning with various task instructions, the text-to-text LLMs, like T5 Raffel et al. (2020), OPT Zhang et al. (2022a) and Llama Touvron et al. (2023), achieved remarkable few/zero-shot generalizations by following these complex instructions Wang et al. (2023b); Ivison et al. (2023).
Trait
NLI-oriented LLM-oriented Human-oriented
Update LLM parameter?
yes maybe yes
Require super large LLMs?
no yes no
Require further label mapping (e.g., verbalizer)?
yes yes no
End-user friendly?
no no yes
Instruction granularity
sentence-level (brief) sentence-level (brief)
paragraph-level (complex)
Instruction scope
output-wise input-wise
task-wise
Task scope
classification classification & generation
classification & generation
Modeling objective
NLI language modeling
follow instructions
Source of indirect supervision
NLI language modeling
various Text-to-Text tasks
@@ 4.2 An Indirect Supervision Perspective
Specifically, NLI-oriented Instructions transform target NLP problems into a source task—NLI—so that the rich supervision from existing NLI datasets can act as indirect supervision for those target problems. LLM-oriented Instructions reformat target problems into the source task—language modeling, so that the rich generic-purpose knowledge in those LLMs can be directly utilized to get the output. Whether it is NLI-oriented Instructions or LLM-oriented Instructions, both try to solve unseen tasks with a generalizable system. However, both of them have limited application scope, e.g., they cannot efficiently deal with some structured prediction tasks Chen et al. (2022b); Zhang et al. (2023). Instead of seeking supervision from a single source task (NLI or language modeling), Human-oriented Instructions learn indirect supervision from a large set of training tasks, the resulting system, therefore, can ideally generalize to any unseen textual tasks. Table 4 further compares them from different dimensions.

### v8

# Large Language Model Instruction Following: A Survey of Progresses and Challenges

@@ Abstract
Task semantics can be expressed by a set of input-output examples or a piece of textual instruction. Conventional machine learning approaches for natural language processing (NLP) mainly rely on the availability of large-scale sets of task-specific examples. Two issues arise: first, collecting task-specific labeled examples does not apply to scenarios where tasks may be too complicated or costly to annotate, or the system is required to handle a new task immediately; second, this is not user-friendly since end-users are probably more willing to provide task description rather than a set of examples before using the system. Therefore, the community is paying increasing interest in a new supervision-seeking paradigm for NLP: learning to follow task instructions, i.e., instruction following. Despite its impressive progress, there are some unsolved research equations that the community struggles with. This survey paper tries to summarize and provide insights to the current research on instruction following, particularly, by answering the following questions: (i) What is task instruction, and what instruction types exist? (ii) How to model instructions? (iii) What are popular instruction following datasets and evaluation metrics? (iv) What factors influence and explain the instructions’ performance? (v) What challenges remain in instruction following? To our knowledge, this is the first comprehensive survey about instruction following. [^1] The curated paper list can be found at:  https://github.com/RenzeLou/awesome-instruction-learning
@@ 1 Introduction
One goal of AI is to build a system that can universally understand and solve new tasks. Labeled examples (Figure 1 (a)), as the mainstream task representation, are costly to obtain at scale or even do not exist in some cases. Then, is there any other task representation that can contribute to task comprehension? Textual instructions provide another dimension of supervision for expressing the task semantics, which often contains more abstract and comprehensive knowledge of the target task than individual labeled examples. As shown in Figure 1 (b), with the availability of task instructions, systems can be quickly built to handle new tasks. Such efficiency is highly desirable in real-world applications, especially when task-specific annotations are scarce. More importantly, instruction following leans toward human intelligence in terms of learning new tasks—a little child can well solve a new mathematical task by learning from its instruction and a few examples Fennema et al. (1996); Carpenter et al. (1996). As a result, this new learning paradigm has recently attracted the main attention of the machine learning and NLP communities Wang et al. (2022b); Longpre et al. (2023).
@@ 2.1 Instruction Following
As illustrated in Figure 1, unlike traditional example-driven supervised learning, the essence of instruction following is to train the LLMs to understand various instructions and produce the corresponding responses. Since this capacity can be extended to any unseen downstream tasks, instruction following has become an efficient learning paradigm for solving few/zero-shot tasks (Radford et al., 2019; Schick and Schütze, 2021c; Yin et al., 2022; Li et al., 2023a; Gupta et al., 2023; Sun et al., 2024; Xie et al., 2024b, inter alia). However, the performance of instruction following highly relies on both model and task scale: a larger LLM (or pretraining with more tokens) tuned on more diverse tasks can achieve significantly better few/zero-shot performances on the downstream tasks (Chung et al., 2022; Iyer et al., 2022; Wang et al., 2023c, inter alia). As scaling model size is unrealistic for most of us, numerous recent studies worked on collecting high-quality instruction-tuning datasets, either employing human labors Khashabi et al. (2020); Ye et al. (2021); Sanh et al. (2022); Wang et al. (2022b); Longpre et al. (2023); Conover et al. (2023); Köpf et al. (2023) or distilling supervision from the powerful LLMs Wang et al. (2023d); Honovich et al. (2023a); Taori et al. (2023); Peng et al. (2023); Xu et al. (2023b); Geng et al. (2023); Chiang et al. (2023); Xu et al. (2023a); Köksal et al. (2023); Kim et al. (2023); Ding et al. (2023); Yin et al. (2023); Lou et al. (2024), e.g., utilizing ChatGPT or GPT-4 to develop creative task instructions OpenAI (2022, 2023).
@@ 2.2 Surveys on In-context Instructions
Several existing surveys Dong et al. (2023); Huang and Chang (2023); Qiao et al. (2023); Yu et al. (2023a) share similar motivations with us while focusing on merely some sub-area of instruction following, such as prompt, few-shot demonstrations, chain-of-thoughts reasoning, etc. For example, Liu et al. (2023a) provided a comprehensive overview of prompt learning and LLMs, where the prompt can be regarded as one specific type of textual instruction (as categorized in § 4). Some other studies surveying “soft instruction”, namely parameter-efficient fine-tuning methods Lialin et al. (2023), also differ from our scope of “textual instruction”. Notably, Zhang et al. (2023b) also proposed a survey on instruction tuning, however, they mostly focused on the existing datasets and models; while we present a more complete and consistent story of the instruction following, including the instruction categories, modeling strategies and so on, which previous works have never introduced. To the best of our knowledge, this is the first work that provides a comprehensive and high-level story of instruction following.
@@ 3 Preliminary
• Template (T): a textual template that either tries to express task intent or is used for bridging X and Y. [^2] A plain template connecting X and Y, e.g., “The input is [$\ldots$] The output is [$\ldots$]”, no task-specific semantics. T may not be an instruction yet.
@@ 4 What is Task Instruction?—A Unified Perspective from Indirect Supervision
{tblr}
width=colspec = X[l,0.09] || X[l,0.36] | X[l,0.44], rowspec = Q[b]Q[m]Q[m]Q[m]Q[m]Q[m], row1 = bg=azure6, fg=white, font=, roweven = gray!15, rowhead = 1, hspan = minimal,
Task & NLI premise (i.e., input text) NLI hypothesis (i.e., instructions Y)
@@
[Donald Trump]ent served as the 45th president of the United States from 2017 to 2021. ($\checkmark$) Donald Trump is a politician
@@
[Donald Trump]ent1 served as the 45th president of the [United States]ent2 from 2017 to 2021. ($\checkmark$) Donald Trump is citizen of United States
@@
In [1997]time, the [company]sub [hired]trigger [John D. Idol]obj to take over Bill Thang as the new chief executive. ($\checkmark$) John D. Idol was hired.
@@
Salesforce and Slack Technologies have [entered]event1 into a definitive agreement] under which Salesforce will [acquire]event2 Slack. ($\checkmark$) Salesforce acquires Slack after it enters into the agreement with Slack Tech.
@@
{tblr}
[…]
Sentiment Analysis I would like to buy it again. [X] The product is    . Great
@@
$\ldots$ Positive
Entity Tagging [Donald Trump]ent served as the 45th president of the United States from 2017 to 2021. The entity in [X] is a    class? Politician
@@
$\ldots$ People
Relation Tagging [Donald Trump]ent1 served as the 45th president of the [United States]ent2 from 2017 to 2021. [X] entity1 is the    of entity2? Executive
@@
$\ldots$ President
Textual Entailment [X1]: Donald Trump served as the 45th president of the United States from 2017 to 2021.
[X2]: Donald Trump is a citizen of United States. [X2]?    , because [X1] Indeed
@@
$\ldots$ Yes
Translation Donald Trump served as the 45th president of the United States from 2017 to 2021. Translate [X] to French:     / été président …
@@ 4.1 Three Types of Instructions
The advantages of NLI-oriented Instructions learning are four-fold: (i) it keeps the label semantics and makes it possible to encode the input-output relations; (ii) it unifies various classification problems into an NLI task; (iii) by making use of the indirect supervision from existing NLI datasets, a model trained on NLI tasks is expected to work on other tasks in a zero-shot manner; (iv) it extends the original close-set indices classification problem into an open-domain label recognition paradigm. Therefore, it has been widely used in a variety of few/zero-shot classification tasks Xu et al. (2023d), such as classifying topics Yin et al. (2019), sentiments Zhong et al. (2021), stances Xu et al. (2022b), entity types Li et al. (2022a), entity relations Murty et al. (2020); Xia et al. (2021); Sainz et al. (2021, 2022), etc.
{tblr}
width=colspec = X[l,0.08] || X[l,0.23] | X[l,0.73] | X[l,0.08], rowspec = Q[b]Q[m]Q[m]Q[m], row1 = bg=azure6, fg=white, font=, roweven = gray!15, rowhead = 1, hspan = minimal,
[…]
@@
 Input: [X]  Output:     Positive
@@
 Input: [X]  Output:     Donald Trump
@@
As shown in Figure 2 (b) and Table 4, the prompt is a representative of the LLM-oriented Instructions, which is usually a brief utterance prepended with the task input (prefix prompt), or a cloze-question template (cloze prompt). It is basically designed for querying the intermedia responses (that can be further converted into the final outputs) from the LLM. Since the prompted input conforms to the pre-training objectives of LLM (e.g., the cloze-style input satisfies the masked language modeling objective Devlin et al. (2019a)), it helps get rid of the reliance on the traditional supervised fine-tuning and greatly alleviates the cost of human annotations. Thus, prompt learning achieved impressive results on a multitude of previous few/zero-shot NLP tasks, like question answering Radford et al. (2019); Lin et al. (2022), machine translation Li et al. (2022c), sentiment analysis Wu and Shi (2022), textual entailment Schick and Schütze (2021a, b), entity recognition Cui et al. (2021); Wang et al. (2022a), etc.
Despite the excellent performance of prompt techniques, there are still two obvious shortcomings with LLM-oriented Instructions in real-world applications. (i) Not User-Friendly. As the prompt is crafted for serving LLMs, it is encouraged to design the prompt in a “model’s language” (e.g., model-preferred incoherent words or internal embedding). However, this LLM-oriented style is hard to be understood by users and often violates human intuitions Gao et al. (2021); Li and Liang (2021); Qin and Eisner (2021); Khashabi et al. (2022). Meanwhile, the performance of prompts highly depends on the laborious prompt engineering Bach et al. (2022), but most end-users are not LLM experts and usually lack sufficient knowledge to tune an effective prompt. (ii) Applications Constraints. The prompt is usually short and simplistic, whereas many tasks cannot be effectively formulated with solely a brief prompt, making prompt hard to deal with the diverse formats of real-world NLP tasks Chen et al. (2022b); Zhang et al. (2023a).
@@
Human-oriented Instructions essentially denotes the instructions used for crowd-sourcing on the human-annotation platforms (e.g., Amazon MTurk). Unlike LLM-oriented Instructions, Human-oriented Instructions  (Figure 2 (c)) are usually some human-readable, descriptive, and paragraph-style information consisting of various components, such as “task title”, “category”, “definition”, and “things to avoid”, etc. (cf. Mishra et al., 2022b) Thus, Human-oriented Instructions are more user-friendly and can be ideally applied to almost any complex NLP task. Table 4.1.1 further shows some representative task examples.
Accordingly, Human-oriented Instructions have attracted much more attention in recent years (Hu et al., 2022b; Gupta et al., 2022; Yin et al., 2022, inter alia). However, due to the complex nature, Human-oriented Instructions are more challenging to encode by vanilla LLMs. For example, off-the-shelf GPT-2 was found to work poorly on following MTurk instructions Wolf et al. (2019); Efrat and Levy (2020). To tame the LLMs better understand the Human-oriented Instructions, follow-up works began to collect large-scale instruction datasets Mishra et al. (2022b); Wang et al. (2022b). All the previous results showed that, after fine-tuning with various task instructions, the text-to-text LLMs, like T5 Raffel et al. (2020), OPT Zhang et al. (2022a) and Llama Touvron et al. (2023), achieved remarkable few/zero-shot generalizations by following these complex instructions Wang et al. (2023c); Ivison et al. (2023b).
{tblr}
width=colspec = X[l,0.34] X[c,0.18] X[c,0.17] X[c,0.20] , rowspec = |[2pt,MyGrey]Q[b] |[1pt,MyGrey,solid]Q[m]Q[m]Q[m]Q[m] |[0.8pt,MyGrey,dashed]Q[m]Q[m]Q[m] Q[m]Q[m]|[2pt,MyGrey] , row1 = font=, rowhead = 1, hspan = minimal, Trait NLI-oriented LLM-oriented Human-oriented
Update LLM parameter? yes maybe yes
Require super large LLMs? no yes no
Require further label mapping (e.g., verbalizer)? yes yes no
End-user friendly? no no yes
Instruction granularity sentence-level (brief) sentence-level (brief) paragraph-level (complex)
Instruction scope output-wise input-wise task-wise
Task scope classification classification & generation classification & generation
Modeling objective NLI language modeling follow instructions
Source of indirect supervision NLI language modeling various Text-to-Text tasks
@@ 4.2 An Indirect Supervision Perspective
Specifically, NLI-oriented Instructions transform target NLP problems into a source task—NLI—so that the rich supervision from existing NLI datasets can act as indirect supervision for those target problems. LLM-oriented Instructions reformat target problems into the source task—language modeling, so that the rich generic-purpose knowledge in those LLMs can be directly utilized to get the output. Whether it is NLI-oriented Instructions or LLM-oriented Instructions, both try to solve unseen tasks with a generalizable system. However, both of them have limited application scope, e.g., they cannot efficiently deal with some structured prediction tasks Chen et al. (2022b); Zhang et al. (2023a). Instead of seeking supervision from a single source task (NLI or language modeling), Human-oriented Instructions learn indirect supervision from a large set of training tasks, the resulting system, therefore, can ideally generalize to any unseen textual tasks. Table 4 further compares them from different dimensions.